### v1

# arXiv:2603.00610v1[cs.SD] 28 Feb 2026

Abstract
While music generation models have evolved to handle complex multimodal inputs mixing text, lyrics, and reference audio, evaluation mechanisms have lagged behind. In this paper, we bridge this critical gap by establishing a comprehensive ecosystem for music reward modeling under Compositional Multimodal Instruction (CMI), where the generated music may be conditioned on text descriptions, lyrics, and audio prompts. We first introduce CMI-Pref-Pseudo, a large-scale preference dataset comprising 110k pseudo-labeled samples, and CMI-Pref, a high-quality, human-annotated corpus tailored for fine-grained alignment tasks. To unify the evaluation landscape, we propose CMI-RewardBench, a unified benchmark that evaluates music reward models on heterogeneous samples across musicality, text–music alignment, and compositional instruction alignment. Leveraging these resources, we develop CMI reward models (CMI-RMs), a parameter-efficient reward model family capable of processing heterogeneous inputs. We evaluate their correlation with human judgments scores on musicality and alignment on CMI-Pref along with previous datasets. Further experiments demonstrate that CMI-RM not only correlates strongly with human judgments, but also enables effective inference-time scaling via top-$k$ filtering. The necessary training data, benchmarks, and reward models are publicly available.  [^1] Code of benchmark and reward model are publicly available at here , and datasets are available at CMI-Pref-Pseudo and CMI-Pref.
Machine Learning, ICML
1 Introduction
The rapid advancement of Artificial Intelligence Generated Content (AIGC) has significantly impacted the creative industries, with music generation emerging as one of the cornerstones for numerous commercial applications in music, movie and entertainment industries (Cao et al., 2025; Ren et al., 2025; Ma et al., 2024). Despite the proliferation of sophisticated generative models, evaluating their outputs remains a fundamental challenge. Unlike text or image generation which sometimes focus on functional or purpose, music evaluation is typically aesthetically subjective and complicated, requiring assessments of musicality and adherence to complex, often multimodal, user instructions.
Developing these evaluation models is hindered by a critical data scarcity. Although large-scale user interaction data exists in music recommendation (e.g., Spotify Million Playlist (Papreja et al., 2019)), it fundamentally captures user-item affinity—a global preference for genre styles or playlists—rather than generative alignment, which demands assessment of perceptual quality and precise instruction-following. Such recommendation datasets lack the fine-grained, comparative rankings of generated samples against complex, multimodal instructions (such as interwoven lyrics, text descriptions, and reference audio) required to train alignment models (Deshmukh et al., 2024).
Consequently, evaluation methodologies have struggled to bridge this gap. Traditional metrics like Fréchet Audio Distance (FAD) (Kilgour et al., 2019) operate at the distribution level, failing to provide the sample-level signals necessary for post-training or filtering. More recent approaches, such as SongEval (Yao et al., 2025), PAM (Deshmukh et al., 2024), and various MOS predictors (Jeong et al., 2025), have advanced the field by offering sample-level scoring. However, these efforts remain fragmented and narrowly specialized. They typically focus on isolated attributes (e.g., only caption alignment) and rely on rigid input assumptions, whereas state-of-the-art music generation models already support flexible input combinations, ranging from simple text prompts to interwoven lyrics and audio references, as illustrated in Figure 1, highlighting a growing mismatch between model capabilities and evaluation methodologies.
We argue that effective evaluation requires compositional alignment, defined here not merely as adherence to simultaneous constraints, but as the capability of a unified model to adaptively agree with human preferences across these optional and varying input conditions – human must determine which of two audio samples is superior in terms of musicality or better aligned with the provided compositional instructions when providing ground truth preference. A framework capable of judging this versatility—handling cases where inputs may be text-only, lyric-guided, or audio-referenced—is currently missing. To bridge this gap, our proposed CMI-RewardBench integrates diverse task-specific datasets to vigorously evaluate whether a single reward model can judge generation quality against the heterogeneous instruction sets inherent to modern AIGC flows.
In this paper, our contributions are three-fold:
  1. 1.
    We construct CMI-Pref-Pseudo, containing 110k samples labeled via a robust pipeline using Qwen3-Omni (Xu et al., 2025a) with consistency filtering. Complementing this, we introduce CMI-Pref, a high-quality corpus of 4,027 pairs annotated by 31 human experts. These annotations capture fine-grained preferences for musicality, alignment, and confidence levels across diverse genres, instruments, and multimodal prompts (including lyrics and audio-to-audio conditioning).
  2. 2.
    We propose CMI-RewardBench, a unified benchmark for music reward models. By integrating existing resources (PAM, MusicEval, Music Arena) with our CMI-Pref test split, this benchmark evaluates models on five distinct tasks ranging from absolute musicality scoring to complex compositional alignment. This unified approach serves as a rigorous testbed for model versatility across optional input settings. Our baseline evaluations on this benchmark expose a significant capability gap, revealing that even state-of-the-art multimodal LLMs (e.g., Gemini-2.5-Pro) struggle to exceed 80% agreement with human preferences.
  3. 3.
    We develop CMI-RM, a family of music reward models supporting compositional conditioning over text, lyrics, and audio. Uniquely supporting all evaluation settings in CMI-RewardBench via a single, parameter-efficient architecture ($\sim$30M), CMI-RM achieves performance comparable to or better than specialized open-source baselines like SongEval. Furthermore, we demonstrate that CMI-RM provides measurable benefits when used for top-k filtering, enabling “inference-time scaling” for music generation.
We release our dataset, benchmark, and model weights to facilitate future research in aligned music generation.
Figure 1: Comparison between traditional music evaluation and the human judge based on compositional multimodal instruction (CMI), which inspire our reward modeling benchmark framework.
2 Related Work
2.1 RLHF for LLMs and MLLMs
Reinforcement Learning from Human Feedback (RLHF) has successfully aligned generative models with human intent by replacing heuristic proxies (e.g., ROUGE) with learned reward models (Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022). This paradigm has successfully extended to multimodal domains, including text-to-image generation (Kirstain et al., 2023; Xu et al., 2023) and video synthesis (Ahn et al., 2024; Liu et al., 2025b), to address aesthetic quality and semantic alignment. Recent research has also applied preference optimization to speech synthesis to improve naturalness and bridge inference-time distribution gaps (Zhang et al., 2024a, 2025b). Concurrently, the use of LLMs as scalable evaluators (Zheng et al., 2023; Gu et al., 2024) has emerged as a robust alternative to human experts. Qhile these works address text, vision, and speech, music evaluation remains under-explored, fragmented. Current music metrics lacking the capacity to judge condition on compositional instructions.
2.2 Evaluation Metrics of Music
Music evaluation typically bifurcated into distribution-level quality assessment and sample-level alignment metrics. However, existing approaches struggle to address the complexity of compositional instructions.
Quality and Subjective Metrics. Distributional metrics like the Fréchet Audio Distance (FAD) (Kilgour et al., 2019) are the standard for global corpus assessment, with recent variants like MAD (Huang et al., 2025b) and KAD (Chung et al., 2025) improving correlation with human perception. For sample-level musicality, MOS predictors such as PAM (Deshmukh et al., 2024), Audiobox (Jeong et al., 2025) and SongEval (Yao et al., 2025) evaluate aesthetic quality. While effective for general audio, high-performing systems like MusicRL (Cideron et al., 2024), WhisQ (Emon et al., 2025), QAMRO (Wang et al., 2025a), and DRAGON (Bai et al., 2025) remain closed-source.
Alignment and LLM Judging. Alignment is primarily measured via contrastive scores like CLAP (Wu et al., 2023), CLaMP3 (Sharma et al., 2024), MuQ-Mulan (Zhu et al., 2025). While music-specific checkpoints improve human-preference alignment (Grötschla et al., 2025), these metrics are largely restricted to text-to-audio pairs and neglect lyrics or audio prompts. Besides, emerging “LLM-as-a-judge” frameworks like AutoMV (Tang et al., 2025) and music recommendation AutoRaters (Chen et al., ) offer scalable, multimodality evaluation with complex instructions. However, these rely on proprietary models and lack an open-source framework for evaluating compositional multiomdal music instructions with lyrics and audio prompts information.
2.3 Preference Dataset and Platform
The development of robust reward models relies on standardized preference datasets and evaluation platforms. Early efforts such as MusicEval (Liu et al., 2025a), SongEval (Yao et al., 2025), and AudioEval (Wang et al., 2025b) provided the first expert-annotated corpora for absolute quality prediction, forming the basis for major community data challenge and benchmarks (Huang et al., 2025a; Ma et al., 2026; Zhang et al., 2025a). To better capture subjective nuances and reduce the cost of data collection, recent work has pivoted toward pairwise preference datasets, such as the AIME dataset (Grötschla et al., 2025). which benchmarks models via large-scale human comparisons .
Inspired by the crowdsourcing success of Chatbot Arena (Chiang et al., ), GenAI Arena (Jiang et al., 2024), and Copilot Arena (Chi et al., ) etc., the recently introduced Music Arena (Kim et al., 2025) provides a live platform for comparative text-to-music evaluation. While these resources represent significant progress, they primarily focus on text-to-music alignment. Our proposed CMI-Pref dataset fills a critical gap by providing the first large-scale preference corpus specifically targeting compositional instructions, including lyrics and audio-to-audio conditioning. And we further propose CMI-RewardBench utilizing most of previous datasets to reduce potential bias on single test set.
3 Method
3.1 Datasets for CMI-RewardBench
To develop and evaluate music reward models under compositional multimodal music instructions, we introduce a comprehensive dataset ecosystem. Besides, we incorporate existing datasets and a new, large-scale multimodal preference corpus to train CMI reward models.
3.1.1 Evaluating Musicality and Text-to-Music Alignment
To establish a foundation for single-modality evaluation, we curate a subset from three established resources:
PAM: We include 500 audio clips from the music subset. Each clip is associated with MOS for musicality and text-music alignment with a text description.
MusicEval: We utilize 413 samples from the test split. These provide expert-validated MOS for musicality (musical impression). Due to the mismatch of audio, filename and text prompt, we omit its MOS for text-music alignment.
Music Arena: We process 2,800 historical interaction datafrom the Arena platform until Dec 2025. To ensure high-quality preference pairs, we remove instances where generation failed and filter out “tie” or “both bad” labels given different user has different tolerance margin when two given audio are similar, yielding 1,340 preference pairs. We categorize Music Arena labels as musicality. Further discussion is shown in Appendix E.3.
3.1.2 CMI-Pref and CMI-Pref-Pseudo: Large-Scale Compositional Multimodal Preference Datasets
We propose a new dataset CMI-Pref in our benchmark to capture human preferences when music is conditioned on CMIs including text descriptions, lyrics, and audio prompts, as well as a large-score dataset CMI-Pref-Pseudo with label provided by LLM-judger.
Table 1: Statistics of CMI-Pref and CMI-Pref-Pseudo compared with previous datasets.
Data set PAM MusicEval SongEval Music Arena CMI-Pref-Pseudo CMI-Pref
Text Condition unavailable
Lyrics Condition unavailable
Audio Condition
#samples 500 2748 2,399 2,800 110k 4,027
#Testing samples 500 413 - 1,340 - 500
#Duration of audio(hours) 0.83 16.62 140.54 88.30 797.34 133.80
#Duration of reference audio(hours) - - - - 123.95 48.56
#Unique Prompts 100 384 - 883 10,213 2,632
#Models & APIs 5 31 7 17 23 23
Data Collection
We distilled audio from a diverse set of 12 models and 11 commercial APIs to ensure a broad distribution of quality and style. For commercial APIs or products, we generated samples using Suno (v3.5, v4, v4.5, v4.5-plus, v5) [^2] https://suno.com/home, Stable Audio 2.0333https://platform.stability.ai/, Minimax-Music-2.0444https://platform.minimaxi.com/, Mureka (v7.5, o2) [^5] https://platform.mureka.ai/, and Loudly [^6] https://www.loudly.com/. These include equal splits of instrumental and vocal tracks (with lyrics) if the model supports lyrics as input, and equal splits of input with and without audio prompts if applicable.
For open-source models, we generated audio from MusicGen (Copet et al., 2023), Stable Audio Open (Evans et al., 2025), YUE (Yuan et al., 2025), SongGen (Liu et al., ), AudioLDM (Liu et al., 2023), AudioLDM 2 (Liu et al., 2024), DiffRhythm(Ning et al., 2025), Levo(Lei et al., 2025), Magenta Lyria-RealTime(Team et al., 2025), Jamify(Liu et al., 2025c), MusicLDM(Chen et al., 2024), and ACE-step(Gong et al., 2025). 35.6% of samples are conditioned on audio prompts for style transfer or continuation in addition to text and lyrics. Audio caption of the audio prompt is provided by Qwen3-Omni as additional text condition if model input cannot support audio prompt.
Following the annotation protocol described in D, 31 annotators constructed CMI-Pref. We then utilized Qwen3-Omni to pseudo-labels to form CMI-Pref-Pseudo.
Annotation and Statistics
CMI-Pref-Pseudo dataset initially include 130k samples, while retaining 110k pairs after consistency checks. It spans 47,546 generations (797.34h) across audio+lyrics (18.3%), audio-only (17.0%), lyrics-only (19.8%), and text-only (44.8%) settings.
Besides, CMI-Pref comprising 4,027 preference samples from generations produced by 23 models, with a total duration of 133.80 hours. The modality distribution remains balanced: audio+lyrics (15.2%), audio-only (20.4%), lyrics-only (14.9%), and text-only (49.5%). We reserved a balanced 500-pair test set and collected 2,574 rationale feedbacks (68% Chinese, 32% English). Table 1 highlights our distinct advantages in scale, duration, and modality diversity over existing benchmarks.
3.2 CMI-RewardBench
We introduce CMI-RewardBench, a unified benchmark designed to evaluate the capability of music reward models in capturing human aesthetic and instructional preferences.
3.2.1 Evaluation Protocol
To ensure robust evaluation across heterogeneous models and datasets, we employ two primary protocols: regression-based correlation and preference-based accuracy. For the PAM and MusicEval datasets, we evaluate reward models on absolute musicality and alignment scoring. Given that different models and datasets utilize varying score ranges, Mean Squared Error (MSE) is insufficient for measuring generalization. Instead, we prioritize relative trend alignment using Linear Correlation Coefficient (LCC) , Spearman Rank Correlation (SRCC) , and Kendall-Tau (K-Tau). For Music Arena and the CMI-Pref test split, we evaluate models on pairwise preference accuracy. Models must determine which of two audio samples is superior in terms of musicality or better aligned with the provided compositional instructions. Accuracy is calculated by comparing model predictions against experts’ annotation.
3.2.2 Baseline Models
We benchmark a diverse set of current state-of-the-art models, categorized by their primary training objective. The musicality baselines includes PAM (Deshmukh et al., 2024), audiobox (Jeong et al., 2025) and SongEval(Yao et al., 2025). PAM is a reference-free metric leveraging audio-language models for general-purpose quality assessment. Audiobox-Aesthetic is an assessment framework providing four specialized metrics: Production Quality (PQ), Production Complexity (PC), Content Enjoyment (CE), and Content Usefulness (CU). SongEval is a model specifically trained on full-length songs using five aesthetic dimensions: coherence, memorability, naturalness, clarity, and overall musicality.
Figure 2: Model architecture of compositional music instruction reward model (CMI-RM).
For text-music alignment, we exam CLAP-Score in both default (Wu et al., 2023) and suggested music-specialized checkpoints (Grötschla et al., 2025), as well as CLAMP3 (Sharma et al., 2024) is a state-of-the-art multimodal alignment model for ABC symbolic notation, waveform and text. MuQ-MuLan(Zhu et al., 2025) is a joint music-text embedding framework utilizing self-supervised music representation learning. Furthermore, we evaluate the zero-shot capabilities of frontier AudioLLMs for all three tasks—musicality, text-music alignment, and compositional instruction alignment. The evaluated suite includes Qwen series and Gemini series: Qwen2-Audio (Chu et al., 2024), Qwen2.5-Omni (Xu et al., 2025b), and Qwen3-Omni, Gemini 2.5 Flash, 2.5 Pro (Comanici et al., 2025), and 3 Pro.
3.3 Compositional Music Reward Modeling
We develop a reward model architecture to handle multimodal conditioning and predict fine-grained human scores.
3.3.1 Model Architecture
Task Formulation.
Given a compositional prompt
$\mathcal{P}=(t,l,a_{\text{ref}}),$
where $t$ denotes optional text description, $l$ optional lyrics, and $a_{\text{ref}}$ optional reference audio, together with an evaluation audio $a_{\text{eval}}$, our reward model predicts human-aligned preferences over generated music along the two complementary dimensions musicality (MUS) and alignment (ALI). The model outputs two scalar scores $(s_{\text{MUS}},s_{\text{ALI}})\in\mathbb{R}^{2}$.
Architecture.
We adopt a two-tower multimodal architecture following Zhang et al. (2024b). One tower encodes the multimodal prompt, while the other processes the target audio. All encoders are frozen and instantiated from MuQ-MuLan. Text descriptions $t$ and lyrics $l$ are encoded separately using the text encoder, while reference audio $a_{\text{ref}}$ and evaluation audio $a_{\text{eval}}$ are encoded using the audio encoder. Empty modalities are treated as zero tensors during training and inference. The encoded prompt components are concatenated and fused using a 4-layer Prompt Transformer:
$\mathbf{h}_{\text{prompt}}=\text{PromptTF}([\mathbf{E}_{t};\mathbf{E}_{l};\mathbf{E}_{a_{\text{ref}}}]),$ (1)
where $\cdot$$\cdot$] denotes sequence concatenation.
To model interactions between the prompt and the generated music, the fused prompt embedding and the evaluation audio embedding are concatenated and processed by a single-layer self-attention Joint Transformer:
$\mathbf{h_{\text{prompt}}};\mathbf{h}_{\text{eval}}=\text{JointTF}([\mathbf{h}_{\text{prompt}};\mathbf{E}_{a_{\text{eval}}}]).$ (2)
We extract the hidden states corresponding to the evaluation audio tokens, apply temporal pooling, and project them through a lightweight MLP to obtain the final scores:
$(s_{\text{ALI}},s_{\text{MUS}})=\text{MLP}(\text{Pool}(\mathbf{h}_{\text{eval}})).$ (3)
3.3.2 Training Strategy
We train the reward model using a two-stage pipeline that leverages both large-scale pseudo-labeled data and high-quality human annotations. Both MUS and ALI heads are optimized jointly throughout both stages. Given a training sample, we compute the musicality-related loss $\mathcal{L}_{\text{MUS}}$ and the alignment-related loss $\mathcal{L}_{\text{ALI}}$ when applicable. The overall training objective is a combination of the two:
$\mathcal{L}_{\text{total}}=0.5\cdot\mathcal{L}_{\text{MUS}}+0.5\cdot\mathcal{L}_{\text{ALI}}.$ (4)
Stage 1: Preference Pre-training.
We first pre-train the model on preference pairs from CMI-Pref-Pseudo. Training is conducted for 2k steps with a batch size of 48. Pairwise preferences are modeled using Bradley–Terry (Bradley & Terry, 1952) formulation. Given a prompt $\mathcal{P}$ and two candidate audio files $A$ and $B$,
$P(A>B)=\sigma\big(s_{\theta}(\mathcal{P},A)-s_{\theta}(\mathcal{P},B)\big),$ (5)
where $s_{\theta}$ denotes the predicted MUS or ALI score depending on the annotation type. The model is optimized using cross-entropy loss, and tied preferences are excluded from training. To mitigate over-confident decision boundaries induced by noisy pseudo labels, we apply label smoothing with a ratio of 0.2 during this stage. Furthur explanation can be found in Appendix C.1.
Stage 2: Expert Fine-tuning.
We then fine-tune the model on a mixture of high-quality human annotations, combining the training split of CMI-Pref and MusicEval, resulting in a total of 6,647 training samples. We use a batch size of 48 and perform early stopping based on validation performance. The selected checkpoint is obtained after 250 optimization steps selected with early stopping. Human annotations appear in two formats: (1) pairwise preferences $(\mathcal{P},A,B)$, trained using Bradley–Terry loss as in Stage 1; and (2) scalar ratings $(\mathcal{P},A,y)$, where $y\in[1,5]$. For scalar ratings, we regress the predicted scores using
$\mathcal{L}_{\text{reg}}=\text{MSE}\big(2\tanh(as+b)+3,\;y\big),$ (6)
where $s$ denotes the predicted MUS or ALI score. The scaling parameters are initialized to $a=0.2$ and $b=0$ during fine-tuning. We drop these constants during inference.
3.3.3 Test-time Scaling
To evaluate the efficacy of test-time scaling (Jin et al., 2025), we conduct experiments using two backbone models: MusicGen-small and Stable-Audio-Open-small. For each of the 2,183 text prompts from the MusicCaps (Agostinelli et al., 2023) dataset eval-split, we generate 10 audio samples (10 sec each) per model. Our reward model serves as a “best-of-N” filter to select the top-performing sample, where $N\in\{1,3,10\}$. We evaluate the effectiveness via subjective A/B testing: Validating whether reward-model selection consistently aligns with human preferences for superior musical quality. The selection criterion is the average of musicality and alignment scores.
4 Discussion
4.1 Benchmark Results
4.1.1 Comprehensive Evaluation on Musicality
Table 2: Musicality results on CMI-RewardBench. For each metric, the best performance is marked in boldface and second with underline.
Musicality PAM (Music Subset) MusicEval (Test Split) Music Arena CMI-Pref
Method&Model LCC SRCC K-Tau LCC SRCC K-Tau ACC ACC
PAM score 0.5873 0.6099 0.4367 0.6466 0.6724 0.4874 63.13% 65.40%
audiobox-CE 0.5283 0.5204 0.3665 0.6393 0.6599 0.4830 64.25% 71.80%
audiobox-CU 0.4645 0.4704 0.3279 0.6272 0.6764 0.4950 67.76% 71.40%
audiobox-PC 0.2505 0.2230 0.1552 0.1225 0.0768 0.0514 58.73% 59.00%
audiobox-PQ 0.4636 0.4513 0.3166 0.6016 0.6335 0.4620 67.54% 73.80%
SongEval-RM 0.6987 0.6977 0.4997 0.7140 0.6949 0.5185 73.88% 72.40%
Qwen2-audio 0.1468 0.1523 0.1120 0.1455 0.2196 0.1585 5.99% 8.60%
Qwen2.5-omni 0.2776 0.2837 0.2144 0.1655 0.1454 0.1145 36.05% 17.40%
Qwen3-omni 0.4155 0.4113 0.3146 0.3693 0.3101 0.2205 59.63% 60.40%
Gemini2.5-flash 0.3813 0.3693 0.2571 0.4188 0.3886 0.2694 64.12% 64.20%
Gemini2.5-pro 0.4463 0.4355 0.3068 0.4966 0.4902 0.3454 69.75% 70.00%
Gemini3-pro 0.5972 0.5967 0.4283 0.6044 0.6018 0.4400 68.85% 65.80%
- w/o f.t.: Distill only 0.4358 0.4304 0.2981 0.5253 0.5117 0.3711 64.25% 70.80%
- w/ f.t.: CMI-Pref 0.6932 0.6988 0.5101 0.7272 0.7315 0.5495 71.41% 77.80%
- w/ f.t.: CMI + MusicEval 0.6367 0.6606 0.4754 0.8195 0.8266 0.6459 73.43% 78.20%
Superior Generalization on Musicality. Table 2 presents a quantitative comparison of our proposed CMI-RM against all baselines. Our method demonstrates robust generalization capabilities across diverse regression tasks. Specifically, the model fine-tuned on CMI-Pref achieves state-of-the-art performance on the PAM music subset (e.g., SRCC of 0.6988). Notably, while specialized baselines such as SongEval-RM exhibit strong performance on the Music Arena platform, their efficacy degrades on the more compositionally complex CMI-Pref dataset (72.40%). In contrast, by incorporating external data (w/ f.t.: CMI + MusicEval), our CMI-RM maintains highly competitive accuracy on Music Arena (73.43%) while achieving state-of-the-art preference alignment on CMI-Pref (78.20%). This indicates that our preference alignment strategy yields robust representations that consistently correlate with granular human ratings across varying evaluation scenarios.
Deficiency of General-Purpose MLLMs. Comparisons with general-purpose MLLMs highlight their limitations in domain-specific sensitivity. Despite the massive parameter scale of models like Gemini 3 Pro and Qwen3-omni, they struggle with fine-grained musical assessment. For instance, on the internal CMI-Pref test set, Gemini 3 Pro and Qwen3-omni achieve accuracies of 65.80% and 60.40%, respectively. Conversely, our CMI-RM achieves 78.20%, demonstrating a substantial margin of improvement. This empirical evidence supports our hypothesis that general instruction tuning alone is insufficient for nuanced aesthetic critique in the audio domain. Furthermore, we observe that the near-zero accuracy of certain AudioLLMs (e.g., Qwen2-audio) primarily stems from poor instruction-following capabilities; these models frequently degenerate into generating music captions or lyrics rather than adhering to the evaluation formatting, resulting in performance significantly below random guess.
4.1.2 Compositional MultiModal Instruction Alignment Evaluation
Table 3: Benchmark Results on Compositional Multimodal Instruction Alignment.
PAM (Music Subset) CMI-Pref (Subsets)
Text-Music CMI-Pref w/o Audio CMI-Pref w/ Audio
Lyrics-Music
Audio-Music
Metrics LCC SRCC K-Tau ACC ACC ACC ACC ACC ACC
CLAP score (default) 0.4692 0.4517 0.3171 60.80% 64.00% - - 62.40% -
CLAP score (music) 0.3192 0.2881 0.1978 67.20% 73.60% - - 70.40% -
MuQ-Mulan 0.4984 0.4741 0.3341 64.80% 68.00% - - 66.40% -
CLAMP3 score 0.2998 0.3013 0.2068 63.2% 62.4% - - 62.8% -
Qwen2-audio -0.024 -0.025 -0.020 0.80% 2.40% 8.80% 14.40% 1.60% 11.60%
Qwen2.5-Omni 0.1529 0.2084 0.1696 31.20% 37.60% 25.60% 32.00% 34.40% 28.80%
Qwen3-Omni 0.5841 0.5907 0.4714 67.20% 60.00% 64.80% 63.20% 63.60% 64.00%
Gemini2.5-Flash 0.3686 0.2454 0.1851 65.60% 56.00% 69.60% 54.40% 60.80% 62.00%
Gemini2.5-Pro 0.4562 0.4179 0.3192 71.20% 63.20% 73.60% 72.00% 67.20% 72.80%
Gemini3-Pro 0.5201 0.5373 0.4047 67.20% 60.80% 68.80% 64.80% 64.00% 66.80%
- w/o f.t.: Distill only 0.3647 0.3547 0.2449 67.20% 72.00% 69.20% 77.20% 69.60% 73.20%
- w/ f.t.: CMI-Pref 0.5200 0.5243 0.3721 64.80% 72.80% 76.00% 82.40% 68.80% 79.20%
- w/ f.t.: CMI + MusicEval 0.4418 0.4321 0.3008 67.60% 72.80% 76.40% 79.20% 70.20% 77.80%
Table 3 displays the compositional alignment evaluation. Unlike standard benchmarks that primarily focus on text-to-music consistency, CMI-RewardBench utilizes the CMI-Pref dataset to evaluate the capability of models to follow complex instructions involving text, lyrics, and reference audio simultaneously.
Comparison with Objective Metrics and General MLLMs. Standard embedding-based metrics and general MLLMs exhibit limitations when processing compositional instructions. For instance, CLAP scores (default) generally yield moderate accuracy ($\sim$60-64% on CMI-Pref Subsets) but inherently lack the architectural support to process audio prompts, resulting in missing evaluations for audio-conditioned tasks. Similarly, while Qwen3-Omni achieves a high correlation on the standard PAM (Music Subset) with an LCC of 0.5841, its performance drops on the more comprehensive CMI-Pref benchmark (e.g., 64.0% ACC on w/ Audio). This discrepancy suggests that standard metrics may oversimplify the alignment problem by omitting lyrics and reference audio constraints. In contrast, our CMI-RM (w/ f.t.: CMI + MusicEval) demonstrates competitive text-centric alignment capabilities, achieving 70.20% accuracy (w/o Audio), while the CMI-Pref fine-tuned variant achieves state-of-the-art performance on audio-centric tasks (w/ Audio) at 79.20%, outperforming frontier models such as Gemini 2.5 Pro (67.2% and 72.8%, respectively). This validates that CMI-RM provides a unified solution for diverse and multimodal generation controls.
Detailed Breakdown by Modality. The breakdown of CMI-Pref Subsets highlights the specific alignment skills imparted by our dataset. A critical finding emerges from the model’s performance on fully compositional conditions. When conditioned on all three modalities concurrently (Text + Lyrics + Audio), the model fine-tuned on CMI-Pref achieves an accuracy of 82.40%. This indicates that CMI-Pref provides essential supervision signals for fine-grained multimodal alignment, a capability that generalist models like Gemini 3 Pro (66.8% on similar constraints) currently lack. Furthermore, for pure textual and lyrics adherence tasks (w/o Audio), the variant w/ f.t.: CMI + MusicEval reaches 70.20% accuracy. This suggests that while CMI-Pref establishes a robust baseline for multimodal handling, the inclusion of MusicEval further refines the model’s sensitivity to semantic and lyrical content.
4.2 The Effectiveness of Different Training Sets
We ablate how training data sources affect CMI-RM under a fixed architecture and identical trainable parameters. All variants share the same two-head setup (MUS/ALI) and differ only in (i) initialization and (ii) fine-tuning data.
Training variants.
Distill only
Pre-trained on CMI-Pref-Pseudo only (pairwise preference, 1.5 epochs).
Distill+CMI-Pref
Initialized from Distill, then fine-tuned on CMI-Pref (train split).
Distill+MusicEval
Initialized from Distill, then fine-tuned on MusicEval (train split).
Distill+Both
Initialized from Distill, then jointly fine-tuned on CMI-Pref + MusicEval.
Scratch+Both
Random initialization, trained on CMI-Pref + MusicEval.
Table 4: Aggregated ablation results with a fixed CMI-RM architecture. Mean SRCC on PAM averages the musicality and text–music alignment regression tasks. Music Arena accuracy uses the combined score $s_{\text{MUS}}{+}s_{\text{ALI}}$ for pairwise prediction. Mean Acc. on CMI-Pref averages musicality and alignment preference accuracy.
Training variant PAM MusicEval Music Arena CMI-Pref
Metric $\uparrow$ Mean SRCC SRCC Acc. Mean Acc.
Distill only 0.3925 0.5117 64.78% 71.10%
Distill+CMI-Pref 0.6116 0.7315 71.12% 75.90%
Distill+MusicEval 0.4338 0.3460 64.78% 69.00%
Distill+Both 0.5464 0.8266 72.46% 76.05%
Scratch+Both 0.2630 0.4986 71.12% 72.15%
Aggregated metric.
To summarize results across heterogeneous benchmarks, we report one aggregated score per dataset in Table 4. For PAM, we average SRCC over the musicality and text–music alignment regression tasks. For MusicEval, we report SRCC on its musicality MOS. For Music Arena, we report preference accuracy, where the model prediction uses the combined score $s_{\text{MUS}}+s_{\text{ALI}}$ to match the platform’s overall preference signal. For CMI-Pref, we report mean accuracy averaged over musicality and alignment preferences. Detailed per-task numbers are provided in Table 2, Table 3, and Table 4.
Findings.
(1) CMI-Pref is the primary driver for cross-benchmark generalization. Fine-tuning on CMI-Pref consistently improves all four aggregated metrics: Distill+CMI yields consistent improvements over the Distill baseline on PAM (0.6116 vs. 0.3925), MusicEval (0.7315 vs. 0.5117), Music Arena (71.12% vs. 64.78%), and CMI-Pref (75.90% vs. 71.10%). This suggests that high-quality human preference data, when coupled with compositional conditions, provides highly transferable supervision signals.
(2) MusicEval provides a complementary signal, but is insufficient alone. While Distill+MusicEval alone does not yield consistent gains across all sets, Distill+Both effectively integrates both datasets. This joint training substantially improves MusicEval correlation (to 0.8266) while pushing Music Arena and CMI-Pref to their peak performances (72.46% and 76.05%, respectively). We observe a mild trade-off on PAM compared to Distill+CMI, indicating a slight objective mismatch between PAM and MusicEval that the joint training mechanism partially mitigates, while more diverse data is required for better generalization.
(3) Distillation initialization is universally beneficial. Comparing Distill+Both with Scratch+Both, we observe that distillation initialization yields consistent gains across all benchmarks. Most notably, it improves performance on Music Arena from 71.12% to 72.46%, and significantly boosts both PAM (0.2630 to 0.5464) and CMI-Pref (72.15% to 76.05%). These results suggest that our large-scale pseudo-label pre-training (Distill) establishes a robust prior that not only mitigates overfitting but also transfers effectively to noisy, out-of-distribution scenarios such as Music Arena, thereby establishing Distill+Both as the optimal training strategy.
4.3 Test-Time Scaling
(a) MusicGen-small
(b) Stable-Audio-Open-small
Figure 3: Pairwise preference matrices for test-time scaling with RM reranking. Each cell reports the percentage of trials in which annotators preferred system AA (row) over system $B$ (column).
Table 5 shows that RM-based best-of-$N$ reranking provides consistent test-time scaling gains across both backbones. The improvements are more pronounced for MusicGen-small, where MUQ-MULAN and AudioBox/SongEval metrics increase monotonically from N=1N{=}1 to N=10N{=}10. For Stable-Audio-Open-small, the gains are smaller and begin to saturate, suggesting diminishing returns when candidate quality is concentrated.
The human preference matrices in Fig. 3 provide two takeaways. First, Ground Truth (GT) is consistently preferred over all reranked outputs for both backbones, indicating that reranking improves generation quality but does not close the gap to real data. Second, preferences among Top-$k$ selections are not strictly monotonic with $N$, and the gains can saturate (e.g., Stable Audio shows limited differences between N=3N{=}3 and N=10N{=}10). Overall, best-of-$N$ reranking is a simple and effective test-time scaling strategy, with diminishing returns beyond moderate $N$.
Table 5: Test-time scaling results of MusicGen-small and Stable-audio-open-small on objective metrics on text-music alignment (MuuQ-Mulam) and musicality (AudioBox, and SongEval).
Model Alignment AudioBox SongEval
[…]
MusicGen 0.298 6.046 6.989 2.143
MusicGen (N=3) 0.323 6.405 7.255 2.213
MusicGen (N=10) 0.339 6.647 7.416 2.273
Stable Audio 0.293 5.567 7.170 2.055
Stable Audio (N=3) 0.301 5.732 7.245 2.078
Stable Audio (N=10) 0.307 5.799 7.290 2.090
5 Conclusion
We introduced CMI-RewardBench, a unified benchmark for evaluating music reward models under Compositional Multimodal Instruction (CMI), where the models handle optional and heterogeneous conditions (text-only, lyric-guided, and audio-referenced) and aligns with human preferences on both musicality and instruction following. To support this evaluation setting, we built two datasets, combining (i) a large-scale pseudo-labeled corpus (CMI-Pref-Pseudo, 110k pairs) constructed via a robust Qwen3-Omni based pipeline with consistency filtering, and (ii) a high-quality expert-annotated set (CMI-Pref, 4k pairs) covering diverse genres, instruments, and multimodal prompts with confidence annotations.
By integrating CMI-Pref with existing resources (e.g., PAM, MusicEval, Music Arena), CMI-RewardBench provides a rigorous testbed spanning five tasks from absolute scoring to pairwise preference, revealing a clear capability gap that even frontier multimodal LLM judges struggle to reach strong agreement with expert preferences in this setting.
We further developed CMI-RM, a parameter-efficient reward model that supports compositional conditioning over text, lyrics, and audio within a single architecture, achieving performance competitive with or exceeding specialized open-source baselines, and providing measurable gains when used for best-of-$N$ reranking as a simple inference-time scaling strategy. We release the dataset, benchmark, and model weights to facilitate future research.
Acknowledgement
Yinghao Ma is a research student at the UKRI Centre for Doctoral Training in Artificial Intelligence and Music, supported by UK Research and Innovation [grant number EP/S022694/1]. Yinghao Ma also acknowledges the support of Google PhD Fellowship. We thank Termeh Taheri on suggestions on back-end development and Christos Plachouras on feedbacks on platform design.
Impact Statement
This work advances the evaluation and alignment of music generation systems by providing a unified benchmark and lightweight reward models that better reflect human preference under compositional multimodal conditions. Our pipeline includes audio generated via commercial APIs; therefore we follow a TOS-aware release policy: only components that are explicitly redistributable are made public, and any restricted parts are shared (if at all) via an application-based access process. We also acknowledge potential copyright concerns from style/melody similarity in generated music and provide a takedown/correction mechanism. Human preference data are collected via a Music-Arena-style platform under informed consent and data minimization (no personal identifiers; limited metadata), with an opt-out/withdrawal mechanism. Given previous datasets included in benchmark are all CC-BY or CC-NC license, our datasets and benchmark are released under CC-BY-NC-SA license and accompanied by a datasheet/data card documenting sources, licenses/TOS compatibility, and the public release fields.
Appendix A Detailed Metadata of Datasets
A.1 Diversity of Prompts
CMI-Pref contains a highly diverse collection of prompts and lyrics that reflect realistic and heterogeneous user instructions for music generation. Across the full dataset, we collect 10,213 unique prompts, while the human-annotated split contains 2,788 unique prompts. Prompt lengths follow a long-tailed distribution: most prompts are concise style or intent descriptors consisting of a few words, while a non-trivial subset includes longer compositional instructions specifying multiple musical attributes. This range captures both minimal user inputs and more detailed requests that require structured reasoning from reward models.
Semantically, the prompts span a wide variety of musical attributes. They cover a broad spectrum of genres, including popular, electronic, rock, jazz, classical, ambient, folk, and orchestral styles, with many prompts combining multiple genre cues. Beyond genre, prompts frequently specify mood, tempo, instrumentation, and production characteristics, introducing orthogonal dimensions of variation. Such compositional prompts prevent reward models from relying on single-keyword correlations and instead encourage holistic assessment of instruction adherence.
The prompt distribution is also linguistically diverse. While English prompts dominate, the dataset includes non-English and mixed-language prompts, reducing reliance on a single linguistic prior and better reflecting global usage patterns. Importantly, prompts are paired with generations produced by a wide range of music generation models and commercial APIs, ensuring that instruction diversity is not confounded with a specific synthesis pipeline.
In addition to text prompts, CMI-Pref includes lyric-conditioned instructions at a meaningful scale. The human-annotated split contains 840 examples with non-empty lyrics, and the full dataset contains 3,896 such examples. Lyrics vary substantially in structure and length, ranging from short repetitive hooks to multi-stanza verses with clear narrative progression. They include both vocal-focused instructions and lyrics intended to be adapted to different musical styles, posing additional alignment challenges beyond text-to-music generation.
Overall, the diversity of prompts and lyrics in CMI-Pref provides a realistic and challenging testbed for reward modeling. By combining short and long instructions, multiple semantic control dimensions, diverse musical styles, and lyric-conditioned inputs, the dataset supports robust evaluation of reward models under compositional and multimodal instruction settings.
A.2 Annotators Agreement of Overlapping Votes
Table 6: Inter-annotator agreement on overlapping votes, computed over 644 vote pairs from 492 comparisons with multiple annotations.
Metric Instruction Following Music Quality
Agreement Count 445 466
Disagreement Count 199 178
Agreement Rate 0.691 0.724
Krippendorff’s $\alpha$ 0.382 0.447
We analyze inter-annotator consistency for comparisons that received multiple independent votes. A total of 492 comparison pairs were annotated by two or more annotators, producing 1,056 votes. For each comparison, we form all pairwise combinations of its votes, resulting in 644 vote pairs.
We measure agreement using agreement rate and Krippendorff’s alpha, computed separately for instruction following and music quality; we adopt Krippendorff’s alpha as it supports agreement estimation when annotator identities are not fixed across vote pairs. As shown in Table 6, both dimensions exhibit relatively high agreement rate. Music quality judgments exhibit slightly higher consistency than instruction following, as they are more intuitive and place lower demands on specialized musical knowledge. After correcting for chance agreement, Krippendorff’s alpha indicates a moderate level of true agreement, which we consider reasonable given the inherent subjectivity of music evaluation.
A.3 Confidence Scores of Human Annotation
We analyze the distribution of annotators’ confidence scores (Fig. 4). Overall, most votes are associated with relatively high confidence. The average confidence for music quality is slightly higher than that for instruction following, consistent with earlier observations that music quality judgments are more intuitive.
Figure 4: Distribution of annotators’ confidence scores for instruction following and music quality.
Table 7: Agreement between instruction-following and music-quality preferences, and the corresponding average confidence.
Metric Value
#Votes with same preference 3309
#Votes with different preference 718
Agreement rate 0.822
Avg. confidence (same Pref.) IF: 3.674, MQ: 3.734
Avg. confidence (diff. Pref.) IF: 3.171, MQ: 3.311
We further examine the relationship between agreement across instruction following and music quality and annotator confidence (Table 7). Overall, the two dimensions show a high agreement rate, and votes have higher average confidence when they agree. This suggests that disagreements between instruction following and music quality often correspond to more ambiguous or trade-off cases, where annotators are less certain about the overall preference.
Alignment among heads: We observed a high consistency between musicality and instruction-music alignment preferences. In pseudo-labels, the agreement is 91% between musicality and instruction alignment. In human data, this agreement is lower (81%), but conflicts typically occur when annotators report high confidence in instruction alignment despite lower musicality, highlighting the necessity of evaluating these dimensions separately.
Appendix B Pseudo-label Generation for CMI-Pseudo
B.1 Label Acquisition Protocol
To ensure the reliability of our pseudo-labels, we address the well-documented phenomenon of positional bias in Large Language Models (LLMs), where the model’s preference is influenced by the presentation order of the options rather than the content quality alone.
Our protocol for collecting pseudo-labels adopts a Position-Consistency strategy. For a given pair of audio samples $(A,B)$ generated from the same prompt $P$, we conduct a bidirectional assessment:
  1. 1.
    Forward Pass: We query the model with the sequence $(A,B)$ to obtain preference $L_{fwd}$.
  2. 2.
    Reverse Pass: We swap the positions to $(B,A)$ and query the model again to obtain preference $L_{rev}$.
A pseudo-label is considered valid and retained only if the judgment is invariant to position—that is, the model prefers the same underlying audio clip in both the forward and reverse passes ($L_{fwd}=L_{rev}$). Comparisons yielding conflicting results or inconsistent ties are discarded as hallucinations or uncertain boundaries.
B.2 Bias Analysis and Dataset Statistics
We initially sampled 129,545 pairwise comparisons from our generated audio pool. As demonstrated in Table 8, we observe significant distributional shifts when the presentation order is swapped, confirming the presence of positional bias. For instance, in the Musicality metric, the win rate for Candidate A fluctuates from 51.96% in the original configuration to 59.27% in the reversed configuration.
By applying our consistency filter, we retain only the high-confidence labels, resulting in 114,694 valid musicality labels and 117,828 valid alignment labels. As shown in the “Agreed” rows of Table 8, the resulting distribution stabilizes, effectively mitigating the variance introduced by the model’s sensitivity to input order. For the final dataset construction, we retained the intersection of these valid subsets to ensure quality across all dimensions, yielding approximately 110k pairs.
Table 8: Distribution of pseudo-labels across varying query configurations. “Original” and “Reversed” denote the presentation order of the audio pair. “Agreed” represents the subset of labels where the model’s preference remained consistent across both permutations. The discrepancy between Original and Reversed highlights positional bias, which is rectified in the Agreed set.
Configuration Win A (%) Win B (%) Tie (%)
Musicality
      Original $(A,B)$ 51.96 40.96 7.08
      Reversed $(B,A)$ 59.27 33.48 7.25
      Agreed (Filtered) 57.80 36.97 5.22
Alignment
      Original $(A,B)$ 33.65 43.71 22.64
      Reversed $(B,A)$ 38.75 38.84 22.77
      Agreed (Filtered) 36.05 41.23 22.73
The “Original” and “Reversed” distributions should theoretically be identical if the evaluator were perfectly unbiased. The observed divergence necessitates our rigorous filtering approach. We did not shuffle the comparison pairs during analysis, therefore its normal that the labels A, B are uneven.
Appendix C Ablation Study on Pseudo-Labeled Data
C.1 Distribution Shift and the Effectiveness of Label Smoothing
We first pretrain the reward head on pseudo-labeled data until convergence ($\sim$14k steps), and then fine-tune it on real preference data. Although pretraining on AI feedback already yields reasonably high accuracy on the real test set (71.0% on Musicality and 71.8% on Alignment), it provides only limited additional benefit after downstream fine-tuning: pseudo-pretraining without smoothing reaches 75.2%/71.0% after fine-tuning, comparable to directly fine-tuning from scratch (75.0%/69.3%).
We attribute this to a distribution shift between pseudo-labeled and real data, which manifests as over-confident probabilities. Specifically, we track the cross-entropy loss on the real test set, We measure binary cross-entropy (CE)
$\mathcal{L}_{\mathrm{CE}}=-\bigl(y\log p_{\theta}(A\succ B)+(1-y)\log(1-p_{\theta}(A\succ B))\bigr),$
and observe that it can exceed 1.2 initially, while a random guess baseline yields $\ln(0.5)=0.693$ for a binary task. This indicates that the model can be confidently wrong under the shift, which is penalized heavily by cross-entropy even when accuracy changes only slightly. Figure 5 shows that without label smoothing, the CE loss on the pseudo validation set remains stable, whereas the CE loss evaluated on real data increases substantially. In contrast, the corresponding accuracy curves in Figure 6 show much smaller differences, suggesting the model produces over-confident decision boundaries.
To mitigate this over-confidence issue, during pseudo training, we apply label smoothing with parameter $\varepsilon$:
$\tilde{y}=(1-\varepsilon)y+\varepsilon/2,$
which replaces hard targets $y\in\{0,1\}$ with softened targets $\tilde{y}\in(0,1)$. We also only train the model 1.5 epochs. In our experiments, we set $\varepsilon=0.2$, i.e., $(1,0)\rightarrow(0.9,0.1)$. This simple change yields clear downstream gains: pseudo-pretraining with smoothing achieves 77.8% on Musicality and 74.0% on Alignment after fine-tuning, outperforming both direct fine-tuning and pseudo-pretraining without smoothing.
Table 9: Real-test reward-head accuracy (%) under different training pipelines.
Training pipeline Musicality Acc. Alignment Acc.
Pseudo pretrain (before FT) 71.0% 71.8%
Pseudo pretrain $\rightarrow$ FT (no smoothing) 75.2% 71.0%
Direct FT (from scratch) 75.0% 69.3%
Pseudo pretrain $\rightarrow$ FT (with smoothing) 77.8% 74.0%
Figure 5: Cross-entropy loss of checkpoints trained on CMI-Pseudo. Pseudo indicates metrics on the CMI-Pseudo validation set, and Real denotes results on the test set of CMI-Pref.
Figure 6: Accuracy of checkpoints trained on CMI-Pseudo. Pseudo indicates metrics on the CMI-Pseudo validation set, and Real denotes results on the test set of CMI-Pref.
C.2 Ablation on Pseudo-Label Data Size
We ablate the amount of pseudo-labeled data used for pre-training. To simplify the evaluation protocol, we fine-tune only on CMI-Pref. Unless otherwise noted, label smoothing is enabled during pseudo pre-training.
Metrics.
We report two groups of metrics. Transfer before fine-tuning: we evaluate the pretrained model on the CMI-Pref test set and report (i) Pref-Test Acc., defined as the average reward-head accuracy over the Musicality and Alignment heads, and (ii) Pref-Test CE, defined as the corresponding average cross-entropy. Downstream after fine-tuning: after fine-tuning on CMI-Pref, we report the head accuracies: Mus-Acc (Musicality head accuracy) and Align-Acc (Alignment head accuracy).
Setup.
The full pseudo-labeled dataset contains 110k examples. We vary the pseudo data size in $\{4\text{k},8\text{k},16\text{k},32\text{k},64\text{k},110\text{k}\}$. We primarily compare models at a fixed epoch progress during pseudo pre-training (Table 10), which provides a natural normalization across different dataset sizes. For completeness, we also include a fixed-step comparison in Table 11.
Results.
Under the same-epoch protocol (Table 10), increasing the pseudo data size generally improves both transfer (Pref-Test Acc/CE) and downstream performance. Moving from small pseudo sets (4k–16k) to 32k–64k yields noticeably stronger results, while performance saturates from 64k to 110k. Notably, 64k achieves performance comparable to using the full pseudo-labeled set under the same-epoch protocol, suggesting diminishing returns beyond this scale. Therefore, we adopt 64k pseudo-labeled examples as a compute-efficient operating point in subsequent experiments. Since directly finetuning on the 3.5k CMI-Pref yields Pref-Text Acc of 0.7215, we validate our rule-of-thumb that Psuedo data should be more than 10 times compared to real data. The ablation on epoch vs step also demonstrates that epoch is a more natural normalization in this finetuning setting.
Table 10: Pseudo data size ablation under the fixed epoch-progress protocol (same-epoch comparison).
[…]
4k 0.582 0.674 0.726 0.664
8k 0.640 0.656 0.750 0.692
16k 0.605 0.655 0.754 0.700
32k 0.646 0.627 0.782 0.700
64k 0.711 0.593 0.778 0.740
110k 0.720 0.559 0.761 0.735
Table 11: Pseudo data size ablation under the fixed-step protocol (same-step comparison).
[…]
4k 0.588 0.724 0.742 0.696
8k 0.623 0.685 0.768 0.694
16k 0.646 0.636 0.760 0.682
32k 0.566 0.725 0.752 0.710
64k 0.711 0.593 0.778 0.740
110k 0.712 0.570 0.762 0.714
Appendix D Human Annotation Details
D.1 Annotation Protocol for CMI-Pref
D.1.1 Core Objectives
The annotation process is decomposed into three distinct components to ensure a multi-dimensional evaluation of the generated music:
  1. 1.
    Preference Label (A/B): A forced-choice selection between two candidates.
  2. 2.
    Confidence Score (1–5): A quantitative measure of the annotator’s certainty, grounded in constraint satisfaction (for alignment) or quality delta (for musicality).
  3. 3.
    Free-text Feedback: Qualitative justifications focusing on fine-grained details that discrete labels cannot capture.
D.1.2 General Principles
  • Instruction-First: Annotators must strictly evaluate the instruction/prompt before listening to avoid post-hoc rationalization.
  • Holistic and Granular Review: Each sample is evaluated for overall coherence as well as specific details (instrumentation, emotion, structure, and audio fidelity).
  • Dimensional Isolation: Annotators are instructed to decouple Alignment (adherence to the prompt) from Musicality (aesthetic quality and production value). A sample may win in alignment while losing in musicality.
D.1.3 Q1: Textual Music Alignment Preference
Annotators identify which sample better follows the elements specified in the instruction, regardless of aesthetic appeal. Key alignment dimensions include:
  • Instrumentation: Specific instruments (e.g., piano solo, guitar riff).
  • Mood/Atmosphere: Emotional valence (e.g., melancholy, upbeat, tense).
  • Genre/Style/Era: Stylistic markers (e.g., Lo-fi, Baroque, 80s synth).
  • Rhythm/Tempo: Temporal characteristics (e.g., driving drum beat, groovy, energetic).
Confidence Calibration (1–5):
  • 5 (Very Certain): Clear binary distinction; one satisfies all key constraints while the other fails or collapses.
  • 3 (Moderate): Default choice; a perceptible lean toward one candidate without overwhelming dominance.
  • 1 (Uncertain): Highly ambiguous prompts or both samples are indistinguishable in their failure/success.
D.1.4 Q2: Musicality Preference
Annotators evaluate which sample sounds more like a “finished, natural, and professional” musical work, independent of the prompt.
  • Key Criteria: Melodic memorability, structural progression, rhythmic stability, and production clarity (lack of distortion/artifacts).
  • Confidence: Reflects the perceived “quality gap” between candidates.
D.1.5 Q3: Feedback Guidelines
Feedback should consist of 1–3 concise sentences focusing on “audible evidence.” Annotators are encouraged to use specific timestamps and avoid vague descriptors.
  • Positive Justification: “Sample A aligns better due to the presence of the specified saxophone; the melody is more distinct.”
  • Negative Evidence: “Sample B suffers from rhythmic instability at 0:20 and harsh high-frequency distortion.”
D.1.6 Glossary for Annotators
Table 12: Taxonomy of musical attributes used in the annotation process.
Dimension Positive Descriptors Negative Descriptors
Melody Catchy, Memorable, Distinct Repetitive, Generic, Wandering
Structure Coherent, Progression, Build-up Disjointed, Abrupt, Random Loops
Rhythm Groovy, Steady, Driving Off-beat, Unstable, Chaotic
Audio Quality Clean mix, Balanced, High-fidelity Harsh, Muddy, Distorted/Clipping
Vocals Natural, Clear articulation Robotic, Slurred, Artifacts
D.2 Annotation Platform
Our annotation platform is illustrated in Figure 7. It provides a unified interface for pairwise audio comparison, confidence scoring, and free-text feedback collection. Detailed usage instructions are provided in the README included in the submitted supplementary materials.
Figure 7: Platform of human annotation.
D.3 Musicality on MusicArena: Decoding User Preference
To understand how musicality and alignment jointly dictate real-world user preferences, we regress these two predicted dimensions against the overall human preference labels collected from MusicArena. Since MusicArena only provides a single, holistic preference label per pair, we evaluate whether users lean more towards aesthetic musicality or strict instructional alignment.
Table 13: Regression analysis of overall human preference on MusicArena using predicted Musicality (Mus) and Alignment (Ali) scores. The results indicate that general user preference is predominantly driven by musicality.
Method Accuracy (%) AUC MSE
Single Metric Threshold (Threshold = 0)
Musicality Diff 73.4 0.8001 0.1884
Alignment Diff 69.7 0.7714 0.2031
Mus + Ali Diff (Equal) 72.6 0.7962 0.1851
Regression Models (5-fold CV)
Logistic Regression 73.1 0.7991 0.1845
SVM-Linear 72.5 0.7990 0.1846
SVM-RBF 72.6 0.7798 0.1875
Table 15 demonstrates that musicality has an overwhelmingly dominant influence on general preference. A simple thresholding of the musicality difference already yields strong performance, while regression models provide only marginal gains (e.g., +0.3% in accuracy) due to the mitigation of fold-level noise.
Appendix E Discussions on Musicality
E.1 The Definition of Contextual Musicality
Musicality in previous benchmarks, such as SongEval (Yao et al., 2025), isolates absolute acoustic quality. We argue that evaluating ”natural intuitive music preference” requires context.
In our framework, we introduce the concept of Contextual Musicality. Musicality extends beyond mere audio fidelity; it serves as a comprehensive sample of user preference implicitly guided by the generation condition. Creativity, aesthetic alignment, and technical execution collectively shape musicality. By processing the prompt, our model effectively captures the user’s implicit profile and aesthetic appreciation level, making the musicality score a context-aware aesthetic judgment rather than a sterile acoustic measurement.
E.2 Does Prompt Information Contribute to Musicality in CMI-Pref?
To empirically validate whether context improves musicality prediction, we conduct an ablation study using our CMI-Pref dataset. We train two variants of our model on CMI-Pref, optimizing only the musicality head. The first variant drops all prompts during both training and inference (relying purely on self-attention over the evaluated audio), while the second variant utilizes regular conditional training.
Table 14: Ablation study on prompt conditions for Musicality prediction on CMI-Pref. We compare models trained and evaluated with and without compositional conditions across different prompt modalities ($n=125$ for each subset).
Prompt Modality w/ Condition Acc (%) w/o Condition Acc (%) $\Delta$
Text + Audio + Lyrics 82.40 68.80 +13.60
Text + Audio w/o Lyrics 79.20 68.80 +10.40
Text + Lyrics w/o Audio 72.00 76.00 -4.00
Text w/o Audio & Lyrics 68.80 67.20 +1.60
Overall ($N=500$) 75.60 70.20 +5.40
As shown in Table 14, incorporating prompt conditions significantly enhances the overall musicality prediction accuracy (from 70.20% to 75.60%). The improvement is particularly pronounced when reference audio is present. For instance, the “Text + Audio + Lyrics” and “Text + Audio w/o Lyrics” modalities observe substantial gains of +13.60% and +10.40%, respectively. This phenomenon occurs because the reference audio—often a user’s favorite track—acts as a strong “taste anchor,” providing the model with explicit information about the user’s aesthetic standards. Furthermore, detailed text prompts signal a higher level of musical expertise, helping the model penalize specific generation flaws that a sophisticated listener would find unacceptable.
Conversely, we observe an anomalous performance drop ($4.00\%$) in the “Text + Lyrics w/o Audio” setting when conditions are introduced. We hypothesize that this is due to a bottleneck in the text encoder’s ability to process raw lyric structures (e.g., repetitive rhymes and stanza formatting) without the grounding of an acoustic anchor, introducing semantic noise that marginally disrupts the model’s alignment with human aesthetic judgments.
E.3 Musicality on MusicArena
To understand how musicality and alignment jointly dictate real-world user preferences, we regress these two predicted dimensions against the overall human preference labels collected from MusicArena. Since MusicArena only provides a single, holistic preference label per pair, we evaluate whether users lean more towards aesthetic musicality or strict instructional alignment.
Table 15: Regression analysis of overall human preference on MusicArena using predicted Musicality (Mus) and Alignment (Ali) scores. The results indicate that general user preference is predominantly driven by musicality.
Method Accuracy (%) AUC MSE
Single Metric Threshold (Threshold = 0)
Musicality Diff 73.4 0.8001 0.1884
Alignment Diff 69.7 0.7714 0.2031
Mus + Ali Diff (Equal) 72.6 0.7962 0.1851
Regression Models (5-fold CV)
Logistic Regression 73.1 0.7991 0.1845
SVM-Linear 72.5 0.7990 0.1846
SVM-RBF 72.6 0.7798 0.1875
Table 15 demonstrates that musicality has an overwhelmingly dominant influence on general preference. A simple thresholding of the musicality difference already yields strong performance, while regression models provide no gains (e.g., -0.3% in accuracy) due to fold-level noise.
This dominance is directly reflected in the fitted model weights:
  • Logistic Regression: $\text{coef}_{\text{mus}}=1.2296$, $\text{coef}_{\text{ali}}=0.1991$, $\text{intercept}=0.0592$.
  • Linear SVM: $\text{coef}_{\text{mus}}=1.2229$, $\text{coef}_{\text{ali}}=0.1861$.
This near-optimal alignment with in-the-wild user behavior echoes our annotation protocols.
Appendix F Reward Model
F.1 Ablation Study on Mapping Functions
We investigate the impact of different mapping functions used to transform the predicted preference scores for MOS regression. We compare four settings:
  • None: No transformation is applied.
  • Tanh: The method adopted in our main experiments, using a Tanh activation to bound the output.
  • Linear: A linear projection without Tanh activation.
  • Ordinal: We employ an ordinal regression objective with learnable classification margins during training. Note that for evaluation, we use the model’s raw scalar output to compute metrics, bypassing the learned margins.
To ensure a fair comparison, all models are initialized from the CMI-Pref-Pseudo pretraining checkpoint (2,000 steps) and finetuned on CMI-Pref and MusicEval. As shown in Table 16, we observe that the choice of mapping function does not yield significant performance differences.
Table 16: Aggregated results on mapping ablations
Datasets PAM MusicEval Music Arena CMI-Pref
Metrics $\uparrow$ Mean SRCC SRCC ACC Mean ACC
None 0.5069 0.4659 73.73% 75.3%
Tanh 0.6024 0.4702 73.8% 74.8%
Linear 0.4595 0.4691 73.58% 74.40%
Oridinal 0.6079 0.4665 72.31% 75.80%
F.2 The Effect of Duration on Model Inference
The variable-length nature of audio makes it important to analyze the effect of duration at inference time. Scoring an entire waveform can be memory-intensive, while restricting inference to short windows may reduce compute but risk missing long-range musical structure and text–music relationships. We evaluate three inference strategies using the same reward model (w/ f.t.: CMI + MusicEval):
  1. 1.
    first 10: Use only the first 10 seconds of audio to infer the Musicality and Alignment scores.
  2. 2.
    mean 10: Split the audio into non-overlapping 10-second chunks (hop size 10 seconds), infer scores for each chunk, and take the average as the final score.
  3. 3.
    first 120: Match the training setting by extracting MuQ embeddings from the first 120 seconds (concatenating four 30-second segments), and then infer scores using the learned weights.
Among these, mean 10 is the most compute-intensive. With a maximum duration of 120 seconds and batched inference, its memory overhead is only slightly larger than first 120, but it requires more forward passes.
Table 17 reports accuracy on CMI-RewardBench subsets across duration bins. We do not report metrics on PAM and MusicEval due to their short duration( $\leq 60$ for MusicEval and $\leq$30 for PAM). On Music Arena, first 120 achieves the best performance in most bins and shows clear gains over first 10, suggesting that short excerpts often under-represent relevant musical content. mean 10 attains comparable accuracy to first 120 in several cases, indicating that aggregating local judgments can approximate longer-context inference. On CMI-Pref, the same trend generally holds but is less pronounced: first 10 remains competitive, especially in shorter-duration bins, implying that some preference decisions can be resolved from early content.
Table 17: Accuracy (%) by audio duration across inference variants. Highest per bin is bold.
Duration (s), Acc (%)
Metric Variant [10,30) [30,60) [60,120) [120,500]
Music Arena n 119 88 531 600
first 10 72.3 53.4 64.2 59.5
mean 10 75.6 48.9 75.7 67.0
first 120 75.6 53.4 81.1 67.3
CMI-Pref Musicality n 196 77 73 154
first 10 77.6 71.4 87.7 76.0
mean 10 78.1 76.6 89.0 74.7
first 120 77.6 76.6 90.4 73.7
CMI-Pref Alignment n 196 77 73 154
first 10 68.4 74.0 83.6 66.9
mean 10 70.4 75.3 80.8 69.5
first 120 72.5 75.3 83.6 70.8
Tables 18 and 19 quantify how predicted scores change across inference strategies. We compare configurations using Linear Correlation Coefficient (LCC) and Root Mean Square Error (RMSE). We also report summary statistics of the reference scores using the standard first 120 to contextualize absolute differences.
Across both Musicality and Alignment, increasing duration is associated with higher RMSE and lower LCC when comparing short-context to long-context predictions, indicating that short excerpts become less predictive of long-context scores as track length grows. When comparing first 10 against first 120, the overall RMSE-to-STD ratios are 0.5575 (Musicality) and 0.6133 (Alignment), suggesting that duration mismatch introduces substantial variation relative to the natural score spread and motivating longer-context inference. Notably, Alignment exhibits lower short-vs-long correlation than Musicality, consistent with alignment requiring broader temporal context. Additionally, the correlation for short time and long time is lower for alignment, indicating that alignment needs longer context.
Finally, while first 10 exhibits noticeably weaker agreement with first 120 (lower LCC and higher RMSE), it still attains reasonably strong performance on several subsets, suggesting that short-context inference can be “good enough” for some downstream uses (e.g., providing a coarse preference signal during post-training). At the same time, the gap to first 120 indicates that longer temporal context remains beneficial, and we leave it to future work to develop reward models that more reliably leverage long-range structure and text–music relationships.
Table 18: Effect of inference duration on predicted Musicality (RMSE & LCC between three configurations). Statistics for first 120 is shown
Pair [10,30) [30,60) [60,120) [120,500] Overall
n = 1311 n = 824 n = 346 n = 1639 n = 4120
RMSE
f10 vs f120 0.3022 0.3960 0.5465 0.7734 0.5687
m10 vs f120 0.2289 0.2067 0.4292 0.5014 0.3751
f10 vs m10 0.2836 0.3583 0.3485 0.4478 0.3758
LCC
f10 vs f120 0.9597 0.9318 0.8234 0.7541 0.8728
m10 vs f120 0.9790 0.9837 0.9292 0.9271 0.9597
f10 vs m10 0.9619 0.9445 0.8811 0.8153 0.9159
Ref (first120) score stats
[…]
[min, max] [-2.31, 4.50]
Table 19: Effect of inference duration on predicted Text-Music Alignment (RMSE & LCC between three configurations). Statistics for first 120 is shown
Pair [10,30) [30,60) [60,120) [120,500] Overall
n = 1311 n = 824 n = 346 n = 1639 n = 4120
RMSE
f10 vs f120 0.2726 0.3104 0.4164 0.6110 0.4539
m10 vs f120 0.2084 0.1777 0.3082 0.3410 0.2727
f10 vs m10 0.2568 0.2922 0.2928 0.4022 0.3311
LCC
f10 vs f120 0.9480 0.9255 0.8329 0.6836 0.8417
m10 vs f120 0.9708 0.9764 0.9309 0.9062 0.9490
f10 vs m10 0.9508 0.9367 0.8828 0.7738 0.8933
Ref (first120) score stats
[…]
[min, max] [-1.76, 3.72]
Appendix G Detailed Analysis of Results
G.1 Performance on Music Arena Subsets
Table 20: Benchmark Musicality ACC Results on Music Arena (Time & Data Type Analysis).
Model Total Time Period (Month) Data Type
(%) Jul-Aug Sep Oct Nov Dec Instru Vocal
PAM score 63.13 69.13 75.41 58.56 53.14 55.27 56.84 68.35
audiobox-CE 64.25 67.79 73.77 63.36 60.14 55.27 64.25 64.25
audiobox-CU 67.76 71.36 77.0 66.78 60.14 60.73 68.04 67.53
audiobox-PC 58.73 63.75 70.49 55.82 54.54 48.00 52.88 63.57
audiobox-PQ 67.54 68.01 74.86 67.47 65.04 63.27 68.04 67.12
SongEval-RM 73.88 73.60 78.69 71.92 64.34 78.18 77.27 71.08
Qwen2-audio 35.99 38.26 37.91 31.23 38.24 34.91 30.66 39.63
Qwen2.5-Omni 36.05 40.04 39.89 38.36 24.48 30.55 29.98 41.23
Qwen3-Omni 59.63 59.06 58.47 59.93 60.84 60.36 62.93 56.89
Gemini2.5-Flash 64.12 60.54 62.84 69.52 61.27 66.55 65.84 62.70
Gemini2.5-Pro 69.75 65.99 64.48 71.88 73.24 75.27 75.78 64.69
Gemini3-Pro 68.85 63.59 60.28 74.31 78.63 72.22 76.03 62.46
- w/o f.t.: Distill only 65.37 67.56 68.3 67.8 68.53 55.64 63.26 67.12
- w/ f.t.: CMI-Pref 71.57 71.36 75.41 72.95 65.04 71.27 73.81 69.71
- w/ f.t.: CMI + MusicEval 73.21 74.27 75.96 74.66 65.73 72.00 74.14 72.44
Tables 20 and 21 reveal two dominant factors that influence reward model performance: temporal distribution shift and annotation confidence.
Table 20 shows a clear temporal drift in Music Arena. Several reference-free metrics, particularly the Audiobox variants, exhibit a noticeable decline in accuracy from early to later months, indicating sensitivity to evolving generation quality and user preference distributions. SongEval-RM demonstrates relatively stronger stability across time, while general-purpose multimodal LLMs show substantial month-to-month variance, suggesting limited robustness for fine-grained musicality judgment. In contrast, our models maintain more consistent performance over time and achieve improved robustness on vocal music, especially when jointly fine-tuned with MusicEval, highlighting the benefit of music-specific supervision.
G.2 Performance on CMI-Pref Subsets
Table 21 analyzes performance on the CMI-Pref test set by stratifying samples according to annotator-reported confidence, which in our data collection protocol explicitly measures the perceived preference margin between two candidates. Unlike traditional pairwise annotation schemes that only record a binary choice, annotators are required to select a preferred sample and additionally indicate how strongly one sample is preferred over the other.
Table 21: Benchmark Musicality ACC Results on CMI-Pref (Ablation Analysis).
Musicality Total Confidence Level Data Type
(500) Conf <3<3 (66) Conf =3=3 (128) Conf >3>3 (306) Instru (250) Vocal (250)
PAM score 65.40% 63.64% 64.84% 66.01% 67.60% 63.20%
audiobox-CE 71.80% 63.64% 72.66% 73.20% 70.80% 72.80%
audiobox-CU 71.40% 66.67% 71.88% 72.22% 68.80% 74.00%
audiobox-PC 59.00% 46.97% 54.69% 63.40% 60.00% 58.00%
audiobox-PQ 73.80% 74.24% 78.13% 71.90% 73.20% 74.40%
SongEval-RM 72.40% 63.64% 71.09% 74.84% 70.80% 74.00%
Qwen2-audio 8.60% 4.54% 7.81% 9.80% 5.20% 12.00%
Qwen2.5-omni 17.40% 15.15% 16.41% 18.40% 15.20% 19.60%
Qwen3-omni 60.40% 54.55% 53.12% 64.70% 63.20% 57.60%
Gemini2.5-flash 64.20% 56.57% 57.94% 71.94% 74.40% 54.00%
Gemini2.5-pro 70.00% 48.49% 69.29% 75.96% 78.80% 61.20%
Gemini3-pro 65.80% 57.94% 66.67% 72.67% 73.20% 58.40%
- w/o f.t.: Distill only 70.80% 53.03% 67.19% 76.14% 65.60% 76.00%
- w/ f.t.: CMI-Pref 77.80% 66.67% 76.56% 80.72% 74.00% 81.60%
- w/ f.t.: CMI + MusicEval 78.10% 66.67% 75.39% 81.70% 75.60% 80.60%
We observe a clear monotonic relationship between confidence and model accuracy across all methods. Higher-confidence comparisons, corresponding to larger perceptual gaps in musical quality or instruction adherence, are consistently easier to predict. Lower-confidence comparisons reflect fine-grained distinctions with smaller margins, which naturally impose a more challenging learning problem.
Importantly, our proposed reward models exhibit the largest performance gains in the high-confidence regime. The CMI-Pref fine-tuned model substantially outperforms all baselines when confidence is high, indicating strong alignment with clear and decisive human preferences. At low confidence levels, performance differences across methods narrow, suggesting that improvements on near-tie comparisons are fundamentally limited by the small preference margins rather than model capacity.
Table 22: Benchmark Text-Music Alignment ACC Results on CMI-Pref(Ablation Analysis).
Text-Music Alignment Total Confidence Level Data Type
(250) Conf <3<3 (47) Conf =3=3 (58) Conf >3>3 (145) Instru (125) Vocal (125)
audiobox-CE 60.00% 45.95% 63.64% 61.90% 56.80% 63.20%
audiobox-CU 59.60% 54.05% 63.64% 59.18% 53.60% 65.60%
audiobox-PC 56.00% 35.13% 59.09% 59.86% 55.20% 56.80%
audiobox-PQ 59.60% 54.05% 65.15% 58.50% 55.20% 64.00%
CLAP score 62.40% 51.35% 59.09% 66.67% 60.80% 64.00%
CLAP music score 70.40% 75.68% 60.61% 73.47% 67.20% 73.60%
MuQ-Mulan 66.40% 67.56% 60.00% 69.39% 64.80% 68.00%
CLAMP3 score 62.80% 64.86% 60.61% 63.27% 63.20% 62.40%
Qwen2-audio 1.60% 2.70% 1.51% 1.36% 0.80% 2.40%
Qwen2.5-Omni 34.40% 27.03% 33.33% 36.73% 31.20% 37.60%
Qwen3-Omni 63.60% 56.76% 63.64% 66.67% 67.20% 60.00%
Gemini2.5-Flash 60.80% 48.65% 60.61% 63.95% 65.60% 56.00%
Gemini2.5-Pro 67.20% 54.05% 66.67% 70.75% 71.20% 63.20%
Gemini3-Pro 64.00% 54.05% 66.67% 65.31% 67.20% 60.80%
- w/o f.t.: Distill only 69.60% 65.96% 70.69% 70.34% 67.20% 72.00%
- w/ f.t.: CMI-Pref 68.80% 55.32% 74.14% 71.03% 64.80% 72.80%
- w/ f.t.: CMI + MusicEval 70.20% 51.06% 68.96% 76.89% 67.60% 72.80%
Table 23: Benchmark Audio-Music Alignment ACC Results on CMI-Pref (Ablation Analysis).
Audio-Music Alignment Total Confidence Level Data Type
(250) Conf <3<3 (42) Conf =3=3 (47) Conf >3>3 (161) Instru (125) Vocal (125)
Qwen2-audio 11.60% 11.90% 4.84% 14.38% 8.80% 14.40%
Qwen2.5-Omni 28.80% 26.19% 25.81% 30.82% 25.60% 32.00%
Qwen3-Omni 64.00% 54.76% 61.29% 67.81% 64.80% 63.20%
Gemini2.5-Flash 62.00% 50.00% 61.29% 65.75% 69.60% 54.40%
Gemini2.5-Pro 72.80% 52.38% 72.58% 78.77% 73.60% 72.00%
Gemini3-Pro 66.80% 61.90% 53.23% 73.97% 68.80% 64.80%
- w/o f.t.: Distill only 73.20% 61.90% 63.83% 78.88% 69.20% 77.20%
- w/ f.t.: CMI-Pref 79.20% 66.67% 72.34% 84.47% 76.00% 82.40%
- w/ f.t.: CMI + MusicEval 77.80% 55.95% 78.72% 83.22% 76.40% 79.20%
Overall, these results demonstrate that explicitly modeling preference strength during data collection provides a principled way to analyze reward model behavior beyond binary accuracy, and highlights the effectiveness of our approach in capturing dominant human preference signals relevant for downstream reranking and selection.
Appendix H LLM-as-a-Judge Prompt Architecture
To leverage frontier multimodal Large Language Models (MLLMs) as evaluators for generated music, we designed a structured message pipeline that seamlessly integrates text descriptions, optional lyrics, optional reference audio, and the generated audio candidates. This ensures the MLLM assesses the candidates based on comprehensive compositional multimodal instructions (CMI) and outputs a rigorously structured JSON response.
H.1 Message Pipeline Construction
The input to the MLLM API is constructed sequentially as a list of messages. This pipeline dynamically adapts to the presence of optional modalities (lyrics, reference audio) and experimental settings (few-shot prompting, chain-of-thought reasoning). The exact message sequence is defined as follows:
  1. 1.
    Task Instruction (System)
    Initializes the evaluator role, defines the evaluation criteria (musicality, instruction following), and specifies the expected JSON schema.
    • Role:system
    • Content: [See Section H.2 for specific System Prompts]
  2. 2.
    Few-Shot Examples (System, Optional)
    If few-shot evaluation is enabled, an instruction and a set of predefined examples are appended to the message sequence. However, we disable this option during distillation, since they did not yield better consistency with human preference labels in our tests and led to a substantial increase in context length.
    • Role:system
    • Content: “You will now be shown several examples of audio comparisons to help you understand how to evaluate the audios.” (Followed by few-shot message pairs)
  3. 3.
    User Instruction Context (User)
    Injects the specific CMI used for generation.
    • Role:user
    • Content:
      • [[Text]] “The user prompt was:<Prompt Text>”
      • [[Text]] (If lyrics provided) “The user provided the following lyrics: <Lyrics Text>”
      • [[Audio]] (If audio provided) <Base64 Encoded Reference Audio>
  4. 4.
    Candidate Audios (User)
    Presents the generated audio samples from the models being evaluated.
    • Role: user
    • Content:
      • [[Text]] “Now, listen to first music file that model 1 generated:”
      • [[Audio]] <Base64 Encoded Model A Audio>
    • Role: user
    • Content:
      • [[Text]] “Now, listen to second music file that model 2 generated:”
      • [[Audio]] <Base64 Encoded Model B Audio>
  5. 5.
    Reasoning Instruction (System, Optional)
    If chain-of-thought rationales are required for the dataset (e.g., CMI-Pref generation), an explicit reasoning instruction is appended.
    • Role: system
    • Content: “Explain your choice, no less than 200 words”
  6. 6.
    Output Formatting Reminder (System, Optional)
    For few-shot settings, a strict reminder of the JSON structure is appended to prevent output drift.
    • Role: system
    • Content: “Now, return your result. Remember, your evaluation format should be in json, [JSON Schema Reminder]”
Implementation Note: All API requests enforce the output format by setting response_format={”type”: ”json_object”} to guarantee parsable results.
H.2 System Prompts
We utilized distinct system prompts depending on the granularity of the evaluation task.
Prompt 1: Overall Preference Evaluation
Used for determining the holistic quality of a generated sample, combining both audio quality and prompt adherence into a single assessment.
You are an experienced critic familiar with music composition. You will have a text input, an optional audio input, both consists as the user prompt. You will then be sent 2 music audios generated by two models. Your task is to select which model’s music is better.
Your answer should be in ‘model_a’, ‘model_b’ or ‘both’, ‘neither’ for your choice.
  • model_a means the first audio is better.
  • model_b means the second audio is better.
  • both means they are equally good and hard to distinguish.
  • neither means both are bad.
Besides this choice, you should also give a score to access the overall quality for each audio on a scale of 1-10, where 1 means terrible and 10 means excellent.
  • Please note that current ai-generated music usually has some limitations, so it’s rare to a high score like 9 or 10.
  • Besides your general impression, scoring should also take basic musical elements into consideration, such as:
    • recording quality (no distortion and scratches)
    • instrument validity (does it sound like the real instrument, where ai commonly struggles with orchestral instruments and traditional instruments)
    • genre consistency (does the music fit the requested genre, where there might a request for a blend)
    • rhythm (stability and creativity)
    • melody (creativity and also stability, is there a clear musical idea)
    • musical structure (repetition and variation, does the music evolve over time if it’s a long piece)
  • Each failure in one of these aspects should deduct points from the total score.
“overall_preference”: ”model_a/model_b/both/neither”,
“score_a”: from 1-10,
“score_b”: from 1-10
Prompt 2: Multi-dimensional Evaluation (Musicality & Instruction Following)
Used to disentangle the evaluation into two distinct dimensions: musical quality and prompt alignment.
You are an experienced critic familiar with music composition. You will have a text input, an optional audio input, both consists as the user prompt. You will then be sent 2 music audios generated by two models.
Your task is to select which model’s music is better from two aspects: which model follows the prompt better (Instruction Following) and which model produces more enjoyable music (Music Quality). For each aspect, your answer should be a choice and two scores.
For both instruction following and music quality, your answer should be in ‘model_a’, ‘model_b’ or ‘both’, ‘neither’ for your choice.
  • model_a means the first audio is better.
  • model_b means the second audio is better.
  • both means they are equally good and hard to distinguish.
  • neither means both are bad.
Besides this choice, you should also give a score on each aspect for each audio on a scale of 1-10, where 1 means terrible and 10 means excellent. You can give float scores like 7.5, the max precision is 1 decimal place.
  • Please note that current ai-generated music usually has some limitations, so it’s rare to a high score like 9 or 10.
  • Besides your general impression, scoring should also take basic musical elements into consideration, such as genre, rhythm, melody, arrangement, timbre, and musical structure.
For music quality, you can consider:
  • recording quality (no distortion and scratches)
  • instrument validity (does it sound like the real instrument, where ai commonly struggles with orchestral instruments and traditional instruments)
  • rhythm (stability and creativity)
  • melody (stability and creativity)
  • arrangement (stability: are the instruments appearing in a logical way, creativity: are there interesting musical ideas)
  • musical structure (repetition and variation, does the music evolve over time if it’s a long piece)
  • Each failure in one of these aspects should deduct points from the score.
For instruction following, you can consider:
  • genre consistency (does the music fit the requested genre, where there might a request for a blend)
  • rhythm consistency (does the rhythm match the prompt, e.g., fast/slow, while maintaining musicality)
  • melody consistency (does the melody style match the prompt, e.g., joyful/sad, while maintaining musicality)
  • arrangement consistency (does the arrangement style match the prompt, e.g., correctly using requested instruments no more or less)
  • Each failure in one of these aspects should deduct points from the score.
Your final answer should be in json format, writing all keys.
“music_quality”: ”model_a/model_b/both/neither”,
“instruction_following”: ”model_a/model_b/both/neither”,
“MQ_score_a”: from 1-10, one decimal place allowed, music quality score for model a,
“MQ_score_b”: from 1-10, one decimal place allowed, music quality score for model b,
“IF_score_a”: from 1-10, one decimal place allowed, instruction following score for model a,
“IF_score_b”: from 1-10, one decimal place allowed, instruction following score for model b
Prompt 3: Rationale Generation Variant
When generating pseudo-labels for distillation and interpretability is required, we replace the final JSON instruction in Prompt 2 with the following to explicitly elicit chain-of-thought rationales:
Finally, provide a detailed explanation of your evaluation, covering the strengths and weaknesses of each audio in relation to the prompt and the characteristics of the models. Your final answer should be in json format, writing all keys.
“music_quality”: ”model_a/model_b/both/neither”,
“instruction_following”: ”model_a/model_b/both/neither”,
“MQ_score_a”: from 1-10, one decimal place allowed, music quality score for model a,
“MQ_score_b”: from 1-10, one decimal place allowed, music quality score for model b,
“IF_score_a”: from 1-10, one decimal place allowed, instruction following score for model a,
“IF_score_b”: from 1-10, one decimal place allowed, instruction following score for model b,
“reason”: ”your detailed explanation of the choice, less than 200 words”
Prompt 4: Musicality Evaluation (Strict JSON Variant)
To ensure strict compliance with the required JSON output format and discrete 1–5 scoring scale, we adopt a simplified prompt for general-purpose AudioLLMs, which may otherwise struggle with complex instruction following.
You are a music comparison judge. Compare Audio A and Audio B, then output ONLY scores.
YOUR ONLY TASK: Give numerical scores (1-5) for each audio and pick a winner.
OUTPUT FORMAT - You MUST output EXACTLY this JSON structure:
{”overall_preference”: ”model_a”, ”score_a”: 3, ”score_b”: 2, ”reason”: ”brief reason”}
MANDATORY RULES:
  • score_a MUST be a number from 1 to 5 (1=worst, 5=best)
  • score_b MUST be a number from 1 to 5 (1=worst, 5=best)
  • overall_preference MUST be one of: ”model_a”, ”model_b”, ”both”, ”neither”
  • You MUST give scores. Scores of 0 are FORBIDDEN.
  • DO NOT output music_genre, music_description, key, tempo, or any music metadata.
  • DO NOT describe what the music sounds like.
  • ONLY output the JSON with scores.
EXAMPLES:
{”overall_preference”: ”model_a”, ”score_a”: 4, ”score_b”: 2, ”reason”: ”A has better quality”}
{”overall_preference”: ”model_b”, ”score_a”: 2, ”score_b”: 5, ”reason”: ”B sounds cleaner”}
{”overall_preference”: ”both”, ”score_a”: 3, ”score_b”: 3, ”reason”: ”similar quality”}
Now compare the two audios and output your scores in JSON format:
Prompt 5: Text-Music Alignment Evaluation (Strict JSON Variant)
Similarly, for evaluating how well the generated audios match the textual prompts under strict formatting constraints, we use:
You are a music comparison judge. Compare how well Audio A and Audio B match the given text prompt.
YOUR ONLY TASK: Give numerical scores (1-5) for how well each audio matches the prompt, then pick a winner.
OUTPUT FORMAT - You MUST output EXACTLY this JSON structure:
{”overall_preference”: ”model_a”, ”score_a”: 3, ”score_b”: 2, ”reason”: ”brief reason”}
MANDATORY RULES:
  • score_a MUST be a number from 1 to 5 (1=worst match, 5=best match)
  • score_b MUST be a number from 1 to 5 (1=worst match, 5=best match)
  • overall_preference MUST be one of: ”model_a”, ”model_b”, ”both”, ”neither”
  • You MUST give scores. Scores of 0 are FORBIDDEN.
  • DO NOT output music_genre, music_description, key, tempo, or any music metadata.
  • DO NOT describe what the music sounds like.
  • ONLY output the JSON with scores.
EXAMPLES:
{“overall_preference”: ”model_a”, ”score_a”: 4, ”score_b”: 2, ”reason”: ”A matches prompt better”}
{“overall_preference”: ”model_b”, ”score_a”: 2, ”score_b”: 5, ”reason”: ”B follows the style”}
Now compare the two audios and output your scores in JSON format:
Prompt 6: Audio-Music Alignment Evaluation (Strict JSON Variant)
For evaluating style transfer and adherence to reference audio conditions, we adapt the strict prompt as follows:
You are a music comparison judge. Compare how well Audio A and Audio B match the reference audio style.
YOUR ONLY TASK: Give numerical scores (1-5) for how well each audio matches the reference, then pick a winner.
OUTPUT FORMAT - You MUST output EXACTLY this JSON structure:
{“overall_preference”: “model_a”, “score_a”: 3, “score_b”: 2, “reason”: “brief reason”}
MANDATORY RULES:
  • score_a MUST be a number from 1 to 5 (1=worst match, 5=best match)
  • score_b MUST be a number from 1 to 5 (1=worst match, 5=best match)
  • overall_preference MUST be one of: “model_a”, “model_b”, “both”, “neither”
  • You MUST give scores. Scores of 0 are FORBIDDEN.
  • DO NOT output music_genre, music_description, key, tempo, or any music metadata.
  • DO NOT describe what the music sounds like.
  • ONLY output the JSON with scores.
EXAMPLES:
{“overall_preference”: “model_a”, “score_a”: 5, “score_b”: 3, “reason”: “A matches reference better”}
{“overall_preference”: “model_b”, “score_a”: 2, “score_b”: 4, “reason”: “B captures the style”}
Now compare the two audios and output your scores in JSON format: